Given that the radiation resistance for a short dipole antenna is 75Ω and the length is 217.6mm, at what frequency would one expect this antenna to radiate?
Answer #1

Hey,

Note: Brother in case of any queries, just comment in box I would be very happy to assist all your queries

R=75

l=217.6=217.6*10^-3 m

R=200*(l/lam)^2

lam=c/f

So,

R=200*(l*f/c)^2

f^2=(R/200)*(c/l)^2 where c is 3*10^8

So, f^2=7.12 *10^17

So, f=8.44*10^8 Hz=0.844 GHZ

Kindly revert for any queries

Thanks.